### v1

@@ 4.1 Model alignment as policy optimization
We adopt a standard sequential decision-making formalism for model alignment, following the RL literature on policy-gradient methods and recent work on LLM post-training. [^3] See, e.g. pg-williams; sutton1999; schulman2016high; kool2019; ouyang2022; ahmadian2024; deepseek-r1; kumar25. In this view, the interaction between an LLM agent and its digital environment is a Markov decision process (MDP) $(\mathcal{S},\mathcal{A},R,P,\mu)$: at each discrete time $t=1,\dots,T$ the agent observes a state $s_{t}\in\mathcal{S}$, samples an action $a_{t}\in\mathcal{A}$ from a stochastic policy $\pi_{\theta}(a_{t}\mid s_{t})$, receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1})$, and the environment transitions to $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. A trajectory (i.e. a token sequence for an LLM) is
@@
that combines correctness of the computed score and adherence to the required output format for regex parsability, evaluated against a gold label $y^{*}$ (surrogated by either $\hat{y}_{\textsf{original}}$ or $\hat{y}_{\textsf{new}}$). In this abstraction, varying $y^{*}$’s value corresponds exactly to changing the reward function $R$ in the MDP.
@@
In summary, under this MDP and GRPO formalism, instantiating the benchmark labels $y^{*}$ by $\hat{y}_{\mathrm{original}}$ or $\hat{y}_{\mathrm{new}}$ shapes the reward assignment $R(\tau)$ in (2), and therefore the advantages $A^{(i)}$ in (5). The rest of the policy-optimization pipeline is held fixed, so any difference in the fine-tuned $\pi_{\theta}$’s test accuracy can be attributed to the choice of labels used to define the alignment objective.
4.2 Controlled Study Setup
@@
Our goal is to isolate the causal effect of label choice on an LLM’s model alignment when RL is used. To this end, we run two GRPO training experiments on the same base model that are identical in all respects except for the labels used to define the training reward. In the original-label arm, training rewards (as assigned by (2)) are computed against MedCalc-Bench’s original labels $\hat{y}_{\textsf{original}}$; in the new-label arm, training rewards are computed against our corrected labels $\hat{y}_{\textsf{new}}$. At test time, the 0-1 accuracy of both models are always graded against $\hat{y}_{\textsf{new}}$, which our Phase 2 study (§3.2.1) found to be more clinically accurate than $\hat{y}_{\textsf{original}}$.
@@
Both RL runs are trained on the same subset of 4,5934{,}593 MedCalc-Bench train instances for which Pipeline 2 produced high-confidence $\hat{y}_{\textsf{new}}$ labels. The held-out test set consists of $887$ instances whose $\hat{y}_{\textsf{new}}$ labels were also obtained from Pipeline 2 and partially audited by physicians. In both arms, the policy LLM sees exactly the same collection of prompts and calculator types; the only difference is whether the training reward’s $y^{*}$ is instantiated using $\hat{y}_{\textsf{original}}$ or $\hat{y}_{\textsf{new}}$. The underlying system and user prompts, including chain-of-thought and answer-format instructions, are held fixed across runs; we provide the full prompt template in the Appendix.
@@
In both arms, the trainable policy $\pi_{\theta}$ and the frozen reference policy $\pi_{\text{ref}}$ are initialized from the same Qwen3-8B base model (qwen3), chosen as a widely used open model for academic post-training. $\pi_{\text{ref}}$, being used for KL regularization, remains fixed throughout training, while $\pi_{\theta}$ is updated iteratively by gradient descent to minimize the GRPO loss in Eq. 6. We use the same updated system prompt (which allows for abstention) and require chain-of-thought style completions with a final scalar answer token; no external tools are used during RL training, and the interaction is purely text-based.
@@
We instantiate the policy objective from §4.1 using the trajectory-level reward $r(\tau)$ in Eq. 2 with $\lambda_{f}=0.1$. For the comparison experiment in §4.3, both runs share the same optimizer and hyperparameters: Adam with learning rate $\text{lr}=10^{-5}$, KL regularization coefficient $\beta=10^{-3}$, and no reward standardization or length normalization to ensure advantage estimation is unbiased (drgrpo). Each GRPO minibatch consists of $256$ distinct $(C,q)$ pairs from the MedCalc training set. We maintain the same random seed so that the minibatch partitions and ordering are identical between the two runs. We also conducted a hyperparameter sensitivity analysis over the learning rates $\{10^{-4},10^{-5},10^{-6}\}$ and KL coefficient $\{10^{-2},10^{-3},10^{-4}\}$.
@@
During training, all trajectories are generated from $\pi_{\theta}$ with a temperature of $1.0$, top-$p=1.0$, and a maximum output length of $1600$ tokens. Moreover, the GRPO group size $G=8$, i.e. $8$ trajectories are sampled i.i.d. from $\pi_{\theta}$ for each $(C,q)$ pair in a minibatch. After training, we evaluate the two policies on the $887$-instance held-out test set, with ths same inference configuraiton as training rollouts. The evaluation metric is the 0-1 test accuracy defined in §3.2: we parse the model’s final scalar score (or abstention token) and compare it against the physician-audited $\hat{y}_{\textsf{new}}$ labels, treating abstentions as correct only when $\hat{y}_{\textsf{new}}$ is NA. All decision rules—prompts, reward shape, decoding, and accuracy metric—are held fixed across the two training runs; only the training labels used inside $r(\tau)$ differ.
@@
Both training runs are conducted on an $8\timesH$100 GPU node with two 56-core Intel 8480C CPUs for 9 hours, using the verl (verl) framework for distributed RL training. The experiment run in Fig. 5 consumes $2\times(8\times 9)=144$ GPU hours; if including the sensitivity analysis, our experiment consumes $144\times(3\times 3)=1296$ GPU hours.
4.3 Comparison Results
Figure 5: Impact of Label Alignment on Model Alignment. Test accuracy dynamics for Qwen3-8B trained via GRPO using our clinically corrected labels (green) versus the original benchmark labels (grey). Shaded bands indicate $\pm 1\sigma$ smoothed over a 10-step window. Improving the reward signal’s factual grounding yields a $8.7\%$ absolute gain in the final moving averages of model test accuracy ($71.4\%$ vs. $62.6\%$).
We present the dynamics of test accuracy during training in Figure 5. Both policies converge after approximately 200 GRPO steps. Averaging the final ten evaluation steps, the policy trained on our corrected labels ($\hat{y}_{\textsf{new}}$) achieves a final accuracy of $71.4\%$, whereas the policy trained on the original benchmark labels ($\hat{y}_{\textsf{original}}$) plateaus at $62.6\%$. This constitutes a substantial $8.7\%$ absolute improvement attributable solely to the quality of the reward signal.
To contextualize this result, we note that a $8.7\%$ performance gap is large relative to recent advances in this domain; for instance, it is comparable to the gains often reported between different alignment algorithms (see Table 2). It is also worth noting that while the label discrepancy between $\hat{y}_{\textsf{original}}$ and $\hat{y}_{\textsf{new}}$ is roughly $32\%$ (see §3.2.1), the downstream performance degradation is smaller ($8.7\%$), suggesting that the RL process displays some resilience to noisy rewards, yet still benefits significantly from cleaner supervision.
Crucially, these results were obtained using standard, default hyperparameters for Qwen3-8B ($\text{lr}=10^{-5}$, $\beta=10^{-3}$, $\lambda_{f}=0.1$), selected a priori without searching for a configuration that favors our method. In addition, the sensitivity analysis we performed over a $3\times 3$ hyperparameter sweep confirms that this performance difference is salient across all nine configurations.
5 Discussion
In this work, we showed that a widely used, mostly LLM-labeled clinical computation benchmark has between one-quarter and one-third of test instances that appear mislabeled, affecting both measured performance and downstream model alignment. Our formal analysis and empirical audits explain how the original data-generating process for MedCalc-Bench’s ground truth causes systematic output errors. The physician-in-the-loop system, instantiated via two automated pipelines, identifies substantially more likely label errors per unit of physician time than uniform sampling, and yields revised labels that physicians judge as more clinically aligned.
Our controlled RL experiment was designed not to claim superior benchmark performance, but to test whether label discrepancies have meaningful downstream consequences or are simply absorbed as noise during training. The substantial accuracy gap we observe (8.7%) rules out the latter hypothesis, demonstrating that benchmark label quality is not merely a measurement detail but a consequential factor in model alignment. If one accepts the evidence from §3 that the corrected labels better reflect clinical ground truth, then it follows that models trained on these labels would be better aligned with physician judgment—but we consider this a secondary implication and leave rigorous validation to future work.
Together, these findings suggest that clinical benchmarks built with partial LLM inputs should be treated as living documents, whose labels co-evolve with both physician feedback and model capabilities. In safety-critical domains, such ongoing maintenance is not optional: benchmarks must be revisited as clinical practice and model capabilities change.
5.1 Limitations
Our study has several limitations. First, the human expert reference used to validate label alignment is high quality but small. Only three physicians from a single healthcare system participated in the blinded review, and they independently computed labels for $50$ contentious test instances. While these labels provide a valuable sanity check that our relabeled $\hat{y}_{\textsf{new}}$’s are much closer to expert judgment than the original $\hat{y}_{\textsf{original}}$’s, a larger panel of physician experts, with the same instances receiving multiple independent labels, would be needed to robustly quantify systematic variability.
Second, our automated pipelines depend on a particular family of LLM verifiers (Gemini-2.5-Pro with search and tools) instantiated with a fixed prompting and tool-use strategy. If Gemini’s failure modes are correlated with those of GPT-4 (used in the original MedCalc-Bench labels) or with the downstream models we wish to evaluate, some systematic error patterns may never be surfaced. Our estimates of the mislabel rate are therefore conditional on this specific verifier.
Third, our automated label verification has incomplete coverage. Pipeline 2 produces high-confidence labels for 85% of MedCalc’s test set and 89% of the sampled train set; the remaining instances where the pipeline fails to reach a supermajority are plausibly the hardest cases, and may be precisely those where physician input matters most. We exclude these deferred instances from our main analyses, meaning that both the estimated mislabel rate and the measured effect of label choice on RL alignment are computed on a more auditor-tractable subset.
Finally, our experiments focus on a narrow but important slice of clinical AI: MDCalc-style risk and severity scores derived from English-language case reports. Real-world EHRs are substantially noisier, often interleaving templated text, missing or conflicting entries, and non-English content. We also instantiate our RL study with a single base model (Qwen3-8B) and one algorithm (GRPO). Although our sensitivity analysis suggests the relative benefit of aligned labels is robust across hyperparameter settings, we do not claim that the magnitude of the effect will transfer unchanged to other architectures or task families.
5.2 Future Work
A central lesson from our study is that the real bottleneck in benchmark maintenance is not computation, but carefully allocated human attention. GPU hours are elastic; physicians’ hours are not. This naturally leads to a design question: given a fixed budget of expert feedback, how should we allocate human evaluation to learn as much as possible about benchmark quality?
Conceptually, there is a latent error structure describing where labels are wrong and which calculators are fragile, and each physician review is a costly observation about that structure. Future work could formalize benchmark maintenance as choosing queries that maximize expected information gain about this latent structure per unit of expert time. Our two-pipeline system can be viewed as a heuristic approximation: LLM auditors concentrate mass on contentious cases, and physicians act as high-leverage adjudicators rather than bulk labelers. The same information-efficient perspective applies upstream to benchmark creation—recent work such as GDPval (gdpval) illustrates how task selection and expert effort allocation shape benchmark design in several domains. Coupling explicit uncertainty models over labels with active-learning or bandit-style policies for expert review is a promising direction we leave to future work.
Acknowledgments
Some of the computing for this project was performed on the Stanford Marlowe GPU cluster (marlowe). We thank Stanford University and Stanford Data Science for providing the computational resources at a subsidized rate that made these research results possible.
6 Appendix

### v2

@@ D.4 Model alignment as policy optimization
We adopt a standard sequential decision-making formalism for model alignment, following the RL literature on policy-gradient methods and recent work on LLM post-training. [^5] See, e.g. [44, 38, 33, 17, 29, 2, 13, 19]. In this view, the interaction between an LLM agent and its digital environment is a Markov decision process (MDP) $(\mathcal{S},\mathcal{A},R,P,\mu)$: at each discrete time $t=1,\dots,T$ the agent observes a state $s_{t}\in\mathcal{S}$, samples an action $a_{t}\in\mathcal{A}$ from a stochastic policy $\pi_{\theta}(a_{t}\mid s_{t})$, receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1})$, and the environment transitions to $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. A trajectory (i.e. a token sequence for an LLM) is
@@
that combines correctness of the computed score and adherence to the required output format for regex parsability, evaluated against a gold label $y^{*}$ (surrogated by either $\hat{y}^{\mathrm{original}}$ or $\hat{y}^{\mathrm{new}}$). In this abstraction, varying $y^{*}$’s value corresponds exactly to changing the reward function $R$ in the MDP.
@@
In summary, under this MDP and GRPO formalism, surrogating the reference label $y^{\star}$ by $\hat{y}^{\mathrm{original}}$ or $\hat{y}^{\mathrm{new}}$ shapes the reward assignment $R(\tau)$ in (5), and therefore the advantages $A^{(i)}$ in (8). The rest of the policy-optimization pipeline is held fixed, so any difference in the fine-tuned $\pi_{\theta}$’s test accuracy can be attributed to the choice of labels used to define the alignment objective.
Appendix E Controlled RL Experiment Details
@@
Our goal is to isolate the causal effect of label choice on an LLM’s model alignment when RL is used. To this end, we run two GRPO training experiments on the same base model that are identical in all respects except for the labels used to define the training reward. In the original-label arm, training rewards (as assigned by (5)) are computed against MedCalc-Bench’s original labels $\hat{y}^{\mathrm{original}}$; in the new-label arm, training rewards are computed against our corrected labels $\hat{y}^{\mathrm{new}}$. At test time, the 0-1 accuracy of both models are always graded against $\hat{y}^{\mathrm{new}}$, which our Phase 2 study (§3.3) found to be more clinically accurate than $\hat{y}^{\mathrm{original}}$.
@@
Both RL runs are trained on the same subset of 4,5934{,}593 MedCalc-Bench train instances for which Pipeline 2 produced high-confidence $\hat{y}^{\mathrm{new}}$ labels. The held-out test set consists of $887$ instances whose $\hat{y}^{\mathrm{new}}$ labels were also obtained from Pipeline 2 and partially audited by physicians. In both arms, the policy LLM sees exactly the same collection of prompts and calculator types; the only difference is whether the training reward’s $y^{*}$ is instantiated using $\hat{y}^{\mathrm{original}}$ or $\hat{y}^{\mathrm{new}}$. The underlying system and user prompts, including chain-of-thought and answer-format instructions, are held fixed across runs; we provide the full prompt template in the Appendix.
@@
In both arms, the trainable policy $\pi_{\theta}$ and the frozen reference policy $\pi_{\text{ref}}$ are initialized from the same Qwen3-8B base model,[31] chosen as a widely used open model for academic post-training. $\pi_{\text{ref}}$, being used for KL regularization, remains fixed throughout training, while $\pi_{\theta}$ is updated iteratively by gradient descent to minimize the GRPO loss in Eq. 9. We use the same updated system prompt (which allows for abstention) and require chain-of-thought style completions with a final scalar answer token; no external tools are used during RL training, and the interaction is purely text-based.
@@
We instantiate the policy objective from §D.4 using the trajectory-level reward $r(\tau)$ in Eq. 5 with $\lambda_{f}=0.1$. For the comparison experiment in §2.6, both runs share the same optimizer and hyperparameters: Adam with learning rate $\text{lr}=10^{-5}$, KL regularization coefficient $\beta=10^{-3}$, and no reward standardization or length normalization to ensure advantage estimation is unbiased.[24] Each GRPO minibatch consists of $256$ distinct $(C,q)$ pairs from the MedCalc training set. We maintain the same random seed so that the minibatch partitions and ordering are identical between the two runs. We also conducted a hyperparameter sensitivity analysis over the learning rates $\{10^{-4},10^{-5},10^{-6}\}$ and KL coefficient $\{10^{-2},10^{-3},10^{-4}\}$.
@@
During training, all trajectories are generated from $\pi_{\theta}$ with a temperature of $1.0$, top-$p=1.0$, and a maximum output length of $1600$ tokens. Moreover, the GRPO group size $G=8$, i.e. $8$ trajectories are sampled i.i.d. from $\pi_{\theta}$ for each $(C,q)$ pair in a minibatch. After training, we evaluate the two policies on the $887$-instance held-out test set, with ths same inference configuraiton as training rollouts. The evaluation metric is the 0-1 test accuracy defined in §3.3: we parse the model’s final scalar score (or abstention token) and compare it against the physician-preferred $\hat{y}^{\mathrm{new}}$ labels, treating abstentions as correct only when $\hat{y}^{\mathrm{new}}$ is NA. All decision rules—prompts, reward shape, decoding, and accuracy metric—are held fixed across the two training runs; only the training labels used inside $r(\tau)$ differ.
@@
Both training runs are conducted on an $8\timesH$100 GPU node with two 56-core Intel 8480C CPUs for 9 hours, using the verl framework for distributed RL training.[35] The experiment run in Fig. 3 consumes $2\times(8\times 9)=144$ GPU hours; if including the sensitivity analysis, our experiment consumes $144\times(3\times 3)=1296$ GPU hours.
Appendix F Two $\mathsf{N/A}$ Instances Found by Phase 1 Audit
In the following pages, we present two MedCalc-Bench test set instances (corresponding to example 1 and 3 in the main text Figure 2) whose ground truth would be better labeled as $\mathsf{N/A}$ (clinically meaningless or unsafe for an LLM to give a numeric score output), due to two respective causes: missing data (required inputs not documented) and population mismatch (calculator applied outside validated scope). These instances are drawn from the 279 test instances (26.6%) that Phase 1 (§2.3) flagged their reference label $\hat{y}^{\mathrm{original}}$ as clinically problematic. The full Phase 1 audit results are available in the supplementary data release.
Case 1  $\triangleright$  Unanswerable Due to Missing Data
Patient Note ($C$)
“A 58-year-old male with dyslipidemia, an eight-year history of T2DM, a family history, his mother, of T2DM, with no known micro- or macrovascular complications, was admitted to the emergency department for malaise, epigastric pain, polyuria, and progressive dyspnea which had begun 10 h ago. He had experienced a 2-kg weight loss over the last few days. His usual medications included aspirin 100 mg q24 h, atorvastatin 40 mg q24 h, and metformin 850 mg q8 h, which had been switched to dapagliflozin 20 days before, due to poor glycemic control, with HbA1c 12% (108 mmol/mol). His vital signs included a heart rate of 122 bpm, respiratory rate 33 rpm, blood pressure 142/70 mmHg, temperature 36.1°C, and body mass index 22.5 kg/m2. On physical examination, somnolence, dry skin and mucous membranes, a Kussmaul breathing pattern, and a capillary refill of 3 sec were observed. Blood tests revealed hemoglobin 17.1 g/dL (13.5–18), leukocytes 19.5 × 103 (4–10 × 103), platelets 296 × 103 (150–450 × 103), glucose 248 mg/dL (60–100), creatinine 0.97 mg/dL (0.67–1.17), sodium 136 mmol/L (135–145), potassium 4.7 mmol/L (3.5–5.5), chloride 101 mmol/L (95–112), phosphate 4.9 mg/dL (2.5–4.5), amylase 70 U/L (10–115), lipase 28 U/L (1–67), pH 6.95 (7.35–7.45), pCO2 23 mmHg (35–45), HCO3 5 mmol/L (22–26), lactate 1.8 mmol/L (0–1.5), urine ketone bodies >150 mg/dL (0–0), CK 112 U/L (1–190), CK-MB 7.3 ng/mL (0.1–5), and troponin I 0.07 ng/mL (0.001–0.05). The electrocardiogram (EKG) showed sinus rhythm with right bundle branch block, and nonspecific repolarization abnormalities. Because of the right bundle branch block was not previously known, a new troponin test was performed six hours later with a peak value of 4.28 ng/mL. Treatment with crystalloids, continuous infusion of intravenous insulin, and administration of potassium and sodium bicarbonate were begun in the emergency room (ER). Due to a poor response over the first two hours, with the persistence of lactic acidosis, the patient was transferred to the intensive care unit (ICU), where more aggressive rehydration with crystalloids was started, without further modifications of the original therapeutic plan. Two days later, the patient was discharged from the ICU to the endocrinology ward. Because of his coronary risk factors and the elevated troponin on admission, a coronary angiography was performed, showing triple-vessel disease. Successful bypass surgery without extracorporeal circulation was performed a few days later, with internal mammary artery grafts to the anterior descendent and marginal obtuse arteries and a saphenous vein graft to the right coronary artery. He was discharged 3 days later on Lantus™ Sanofi-aventis S.p.a Valcanello,03012 Anagni (FR), Italia (insulin glargine) 20 IU and Insulina Novorapid™: Novo Nordisk A/S. Hallas Allé, DK-4400. Kalundborg, Dinamarca (insulin aspart) 6-4-4-0 IU subcutaneous insulin, aspirin, clopidogrel, enalapril, bisoprolol, atorvastatin, and furosemide. The oral antidiabetic treatment with dapagliflozin was not restarted. During follow-up by endocrinology, C-peptide, anti-GAD, and IA-2 antibodies were required due to normal BMI and lack of family history suggested T1DM or a latent autoimmune diabetes of adults (LADA) rather than T2DM as etiology of his diabetes. C-peptide was 1 mg/dL (0.9–7.1) with glycemia 214 mg/dL (60–100), anti-GAD <5 U/mL (0–12.5), and anti-IA-2 <7.5 U/mL (0–7.5).”
Score Question ($q$): Albumin Corrected Anion Gap (in mEq/L) Original Reference Label ($\hat{y}^{\mathrm{original}}$): -2.75 Original Label Metadata: {‘Chloride’: [101.0, ‘mmol/L’], ‘Albumin’: [17.1, ‘g/dL’], ‘Bicarbonate’: [5.0, ‘mmol/L’], ‘Sodium’: [136.0, ‘mmol/L’]} Phase 1 AI Auditor Agent’s Analysis (Excerpt)
[…]
Case 2  $\triangleright$  Population Mismatch
Patient Note ($C$)
“A 55-year-old woman presented with a past medical history of smoking until five years ago (60 pack/year), arterial hypertension, chronic obstructive pulmonary disease with no supplementary oxygen use, and chronic hepatitis B infection with no history suggestive of cirrhosis. Her chief complaint was one year of abdominal discomfort associated with an epigastric mass and weight loss of 30 kg. Simple and contrasted abdominal computed tomography (CT) showed hepatomegaly with focal lesions, the largest being 15 cm in segment VIII, and satellite lesions of up to 5.5 cm in segment VII and 3.5 cm in segment III, with similar morphologic characteristics. Abdominal magnetic resonance imaging (MRI) showed a liver with no signs of chronic disease, and an infiltrative dominant lesion of 11 x 8 cm in segments IV and V, associated with smaller lesions in both lobules of 1.6 cm in segment III, and 0.9 cm in segments VII and VIII (Figure ). Subsequently, the patient presented aggressiveness and disorientation, where hypoglycemia was documented (33 mg/dL) and treated with dextrose 10% with complete symptom resolution. Physical examination revealed cachectic facies, and an abdominal mass of approximately 13 cm located in epigastrium and right hypochondrium, associated with hepatomegaly. Initial laboratory were: white blood count of 9900/uL, neutrophils 8000/uL, lymphocytes 1700/uL, monocytes 300/uL, red blood cells 5,540,000/uL, hemoglobin 16.6 g/dL, hematocrit 49.6%, mean corpuscular volume 89.6 fL, platelet count 357,000/uL, international normalised ratio (INR) 1.05 (normal range (NR) 0.8-1.4), partial thromboplastin time 33 seconds (NR 25-35), serum potassium 3.3 mmol/L, chloride 104 mmol/L (NR 98-107), sodium 141 mmol/L (NR 136-145), total bilirubin 1.01 mg/dL (NR 0.2-0.9), direct bilirubin 0.48 mg/dL (NR 0-0.3), alkaline phosphatase 404 U/L (NR 38-110), gamma-glutamyl transferase (GGT) 1017.6 U/L (NR 10-38), lactic dehydrogenase (LDH) 293 U/L (NR 120-246), alanine transaminase (ALT) 124 U/L (NR 9-52), aspartate transaminase (AST) 226 U/L (NR 14-36), serum creatinine 0.22 mg/dL (NR 0.5-0.8), serum alpha fetoprotein (AFP) 23184.1 ng/mL (NR 0.8-1), carcinoembryonic antigen 0 ng/mL, CA 19-9 antigen 26.86 U/mL (NR 0-37), cortisol 8.9 ug/dL (NR 5.27-22.45), blood ketones 1.0 mg/dL (NR < 3.5), serum glucose 17 mg/dL (NR 74-106), baseline insulin 0.40 uU/mL (NR 2.6-24.9), IGF-I < 15 ng/mL (NR 45-210), C peptide 0.0407 ng/mL (NR 1.1-4.4), IgM antibodies against hepatitis B core antigen non-reactive, antibodies against hepatitis B surface antigen < 3.10 mUI/mL (NR 0-10), antibodies against hepatitis C non-reactive, hepatitis B surface antigen (HBsAg) > 1000 (NR 0-1). Given our patient’s profile, a “big” IGF-II measurement was attempted, but the assay is not available in Colombia”……
Score Question ($q$): Sodium Correction for Hyperglycemia (in mEq/L) Original Reference Label ($\hat{y}^{\mathrm{original}}$): 139.392 Original Label Metadata: {‘Glucose’: [33.0, ‘mg/dL’], ‘Sodium’: [141.0, ‘mmol/L’]} Phase 1 AI Auditor Agent’s Analysis (Excerpt)
[…]
Appendix G MedCalc-Bench Usage in Anthropic Press Release
In Anthropic’s “Claude for Healthcare” press release,[3] referenced earlier in the main text’s Introduction section (§1), the company did not provide the evaluation source code or the version number of MedCalc-Bench dataset with which their models obtained the reported accuracies, i.e., whether it was the original peer-reviewed version published in NeurIPS 2024 (v1.0), or a recent GitHub/HuggingFace revision released by the same authors without a technical report on how their reference labels’ curation methodology changed (v1.2). [^6] Appendix §B provides more context on the distinction between the two MedCalc-Bench versions. Identifying the dataset version used by Anthropic is essential for gauging the clinical validity of their LLMs’ benchmark results. To balance inference cost with experiment relevance, we tried to independently evaluate Anthropic’s Haiku 4.5| and
G.1 Reproducing Anthropic’s Results on MedCalc-Bench
This experiment has two treatment groups, one evaluating the two said Claude models on the test set of MedCalc-Bench v1.0 (1047 instances), and the other evaluating the two Claude models on the test set of MedCalc-Bench v1.2 (1100 instances). We set the inference hyperparameters according to their press release’s two footnotes (see bottom of Figure 7), which says “Claude 4.5 models evaluated with extended thinking (64k tokens) and native tool use” and “(with Python code execution)” for MedCalc-Bench. Figure 9 shows the prompt template we used for all API calls.
G.2 Our Experiment Results and Analysis
Figure 8 compares Anthropic’s officially reported MedCalc-Bench accuracy with our independent reproduction attempts using both dataset versions. On MedCalc-Bench v1.0, we obtained 50.7% accuracy for Haiku 4.5 and 62.6% for Sonnet 4.5 (average of two runs: 62.7% and 62.5%). These results are closer to Anthropic’s reported values of 47.9% and 54.4% than our v1.2 results, which were substantially higher at 65.3% and 80.5%, respectively.
The proximity of our v1.0 results to Anthropic’s official figures suggests that the company likely evaluated their models on the original NeurIPS 2024 dataset rather than the later GitHub revision. However, a modest discrepancy remains: our v1.0 reproduction exceeds their reported accuracy by 2.8 percentage points for Haiku 4.5 and 8.2 percentage points for Sonnet 4.5. This gap may reflect differences in prompt formatting, answer extraction logic, or other undisclosed evaluation details.
As a sensitivity analysis aside from the experiment treatment, we also tested whether turning on the web search tool, which Anthropic’s press release left unmentioned its usage, affected performance. On v1.0, adding web search improved Haiku 4.5 accuracy from 50.7% to 53.3% (+2.6pp) but slightly decreased Sonnet 4.5 accuracy from 62.6% to 61.2% (-1.4pp). On v1.2, turning on web search yielded gains of +4.3pp for Haiku 4.5 (65.3% to 69.6%) and +1.6pp for Sonnet 4.5 (80.5% to 82.1%). These modest improvements suggest that web search provides supplementary medical knowledge grounding but does not account for the large version-dependent accuracy differences observed.
Figure 7: Benchmark performance of four Claude-series models was reported by Anthropic in Exhibit #1 of their Jan 11, 2026 press release.[3] Screenshot taken on Jan 17, 2026. We configured our API call pipeline to match the setup described in the first two bullet-point footnotes. MedAgentBench [15] is unrelated to our study.
Figure 8: MedCalc-Bench accuracy comparison between Anthropic’s official results (shown in Figure 7) and our independent reproduction using both dataset versions. Our v1.0 reproduction (50.7% for Haiku 4.5; 62.6% for Sonnet 4.5) better matches Anthropic’s official values (47.9%; 54.4%), suggesting they likely evaluated on the original NeurIPS 2024 dataset. Results on v1.2 are substantially higher, indicating that dataset version has a large effect on reported accuracy. All evaluations used extended thinking (64k tokens) and Python code execution, matching the configuration described in Anthropic’s press release.
Claude API System Prompt: (LEFT EMPTY) Claude API User Prompt Template: You are a medical professional solving a clinical calculation problem. **Patient Note:** {patient_note} **Question:** {question} **Instructions:** 1. Read the patient note carefully and extract all relevant clinical values. 2. Identify the appropriate medical formula or calculation method needed. 3. Use Python code to perform the calculation accurately. Show your work step by step. 4. After completing your calculation, provide your final numerical answer. 5. Your final answer MUST be enclosed in <answer></answer> tags. For example, if your calculated answer is 42.5, you would write: <answer>42.5</answer> Now solve the problem:
Figure 9: API call prompt template, applied identically to both treatment groups in §G.1.